### v1

# Fragmented eigenstate thermalization versus robust integrability in long-range models

Abstract
Understanding the stability of integrability in many-body quantum systems is key to controlling their dynamics and predicting thermalization. While much is known about how integrability breaks down in short-range interacting systems, the corresponding picture for long-range couplings remains incomplete. Yet long-range interactions are both ubiquitous in nature and readily engineered in modern experimental platforms. Here, we show that integrability in fully connected models is either robust or extremely fragile depending on whether the perturbation is non-extensive, extensive one-body, or extensive two-body. In a finite system with short-range interactions, any of these perturbations can induce chaos when applied with finite strength. In contrast, in fully connected finite models, chaos is induced by extensive two-body perturbations, and they do so even at infinitesimal strength. In this case, chaos emerges within quasi-symmetry sectors, leading to a fragmented manifestation of the eigenstate thermalization hypothesis (ETH). This challenges previous claims of ETH violation in quantum systems with strong long-range interactions.
Long-range interacting systems have raised fundamental challenges since their earliest studies, as they violate key assumptions of conventional statistical mechanics. Unlike short-range systems, where additivity and extensivity ensure well-defined thermodynamic limits, long-range interactions induce strong correlations across the entire system, requiring a refined theoretical framework. The conditions for thermodynamic stability in such systems were rigorously established in [1, 2]. Building on this foundation, subsequent studies uncovered other unique features, that include ensemble inequivalence [3, 4], quasi-stationary states [5, 6, 7, 8], and anomalously slow relaxation [9, 7, 8, 10].
In the quantum domain, long-range interacting systems have revealed a variety of nontrivial phenomena rooted in their intrinsic non-locality and non-additivity [11]. These include excited-state quantum phase transitions [12, 13, 14], finite-energy phase transition in one dimension [15], slow entanglement growth [16, 17, 18], violations of Lieb-Robinson bounds [19, 20, 21, 22], cooperative shielding [23, 24], discrete time crystal phases [25, 26], strong prethermalization [27, 28, 29, 18, 30], unconventional dynamical phase transitions [31, 32, 33, 34, 35, 36], and the emergence of robust many-body quantum scars [37]. Motivated by these theoretical predictions and enabled by experimental advances in controlling long-range couplings and achieving long coherence times, particularly in arrays of trapped ions [38, 39, 40, 41] and Rydberg atoms [42, 43], a growing number of experiments have begun to explore far-from-equilibrium dynamics in long-range interacting systems [44]. They have led to direct observations of light-cone violation [40, 41], entanglement generation [45], dynamical phase transitions [46], and long-lived prethermal states [47].
A central open question in systems with long-range interactions concerns the interplay between collective dynamics, prethermalization, and eventual thermalization. Within the framework of the eigenstate thermalization hypothesis (ETH) [48], thermalization in isolated quantum systems arises from the onset of many-body quantum chaos [49]. In systems with short-range interactions, integrability is typically fragile and the addition of even a single impurity can induce chaos [50, 51, 52, 53, 54, 55]. This fragility implies that integrability-breaking perturbations drive the system toward chaos at a strength that decays exponentially with system size. Whether this paradigm extends to systems with long-range interactions remains an open question. In [56], an arbitrarily small perturbation was shown to induce chaos in the presence of long-range interactions, although in [57] a finite perturbation threshold was required for thermalization.
To investigate how many-body quantum chaos emerge, or fail to emerge, and whether ETH holds in systems governed by long-range interactions, we consider a prototypical Hamiltonian, experimentally realized in systems with trapped ions, where the interaction strength decays algebraically with distance as $1/r^{\alpha}$. In the limit $\alpha=0$, the model becomes fully connected and mean-field integrable, featuring a highly degenerate spectrum fragmented into distinct energy bands. Contrary to its short-range counterparts, where any finite integrability-breaking term typically induces chaos, we find that the $\alpha=0$ integrable point is either remarkably robust or exceptionally fragile, depending on the nature of the perturbation. For some classes, integrability persists under finite perturbations, while for others, chaos emerges even at infinitesimal strength. In this fragile case, chaos arises within the fragmented energy bands and ETH becomes satisfied sector by sector. Our results demonstrate that the route to chaos in long-range systems depends sensitively on the nature of the perturbation, challenging conventional expectations based on short-range models.
Model and spectrum.– We consider the experimentally realized [40] one-dimensional system with open boundary conditions described by the following spin-1/2 Hamiltonian with $L$ sites,
$\hat{H}=\mathcal{N}_{\alpha}\sum_{i>j=1}^{L}J\frac{\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{x}}{|i-j|^{\alpha}}+h\sum_{i=1}^{L}\hat{\sigma}_{i}^{z},$ (1)
where $\hat{\sigma}_{i}^{x,z}$ are Pauli matrices at site $i$, and $\mathcal{N}_{\alpha}=1/L^{(1-\alpha})$ for $\alpha<1$ is the thermodynamic Kac’s scaling that ensures energy extensivity for $\alpha\leq 1$. Throughout the paper with fix $J=1$ and $h=1$. The eigenvalues and eigenstates of $\hat{H}$ are denoted by $E_{n}$ and $|n\rangle$.
The model is integrable in both limits: for nearest-neighbor interaction ($\alpha\to\infty$), when it becomes the transverse-field Ising model, and for all-to-all couplings ($\alpha=0$), which has also been experimentally realized [58]. In this case, Hamiltonian (1) becomes equivalent to that of the Lipkin-Meshkov-Glick (LMG) model,
$\hat{H}^{\alpha=0}=\frac{2J}{N}\hat{S}_{x}^{2}+2h\hat{S}_{z}-2J,$ (2)
where the collective spin operators are defined as $\hat{S}_{x,y,z}=(1/2)\sum_{i=1}^{L}\hat{\sigma}_{i}^{x,y,z}$. The LMG Hamiltonian $\hat{H}^{\alpha=0}$ has $SU(2)$ symmetry, as the total spin $\hat{S}^{2}=\hat{S}_{x}^{2}+\hat{S}_{y}^{2}+\hat{S}_{z}^{2}$ is conserved.
The spectrum of $\hat{H}^{\alpha=0}$ is fragmented into highly degenerate energy bands characterized by the total spin quantum number $s$ corresponding to the eigenvalue of $\hat{S}^{2}$. For even $L$, $s$ ranges from 0 to $L/2$, resulting in $(L/2+1)^{2}$ distinct bands [see the Supplemental Material (SM) for further details]. The band structure shown in the density of states (DOS) in Fig. 1(a).
To systematically assess the stability of integrability in $\hat{H}^{\alpha=0}$, we classify perturbations into three experimentally relevant categories. The energy-band structure of the DOS in Fig. 1(a) remains unchanged under an infinitesimal perturbation $\hat{V}$ of strength $\delta$ from any of the three classes. They are schematically depicted in Fig. 1(b) and categorized as follows: (i) Non-extensive perturbations include one-body or two-body terms, local or non-local, whose support does not scale with system size, such as $\hat{\sigma}_{L/2}^{z}$, $\hat{\sigma}_{L/2}^{z}\hat{\sigma}_{L/2+1}^{z}$, or $\hat{\sigma}_{1}^{z}\hat{\sigma}_{L}^{x}$. (ii) Extensive one-body perturbations consist of uniform or disordered fields with support growing with system size, such as $\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}$, $\sum_{i=1}^{L/2}\hat{\sigma}_{i}^{z}$, or $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}$ where $h_{i}\in[-\delta,\delta]$ are random numbers. (iii) Extensive two-body perturbations include local two-body terms with system-size-dependent support, such as $\sum_{i=1}^{L}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}$, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x}$ or infinitesimal changes to the power-law exponent, $\alpha\to\alpha+\delta\alpha$.
Figure 1: In the all-to-all coupling limit ($\alpha=0$), (a) the density of states (DOS) is fragmented into energy bands when the system ($L=8$) is subjected to (b) perturbations from any of the three categories. However, the (c) level statistics within the most populated energy band reveal that only class (iii) perturbations induce many-body quantum chaos, as indicated by the Wigner-Dyson level spacing distribution for system ($L=14$). Parity and inversion symmetries were taken into account.
Despite sharing equivalent DOS, the three classes of infinitesimal perturbations yield markedly different spectral properties. The presence of correlated eigenvalues, as in random matrix theory [59], is a widely used diagnostic of quantum chaos in isolated systems. In particular, short-range spectral correlations can be quantified through the analysis of adjacent energy levels, for example via the distribution of the ratio $r_{n}=\min{({\cal S}_{n},{\cal S}_{n-1})}/\max{({\cal S}_{n},{\cal S}_{n-1})}$ of consecutive level spacings, ${\cal S}_{n}=E_{n+1}-E_{n}$, [60]. Quantum chaos is signaled by level repulsion and Wigner-Dyson statistics [61], whereas integrable systems exhibit uncorrelated levels, with Poissonian spacing distributions, or degeneracies arising from conserved quantities. In Fig. 1(c), we show the level spacing distributions computed within the most populated energy band. Integrability is preserved under class (i) and class (ii) with uniform perturbations – even at finite strength. Class (i) perturbations barely lift degeneracies, while class (ii) uniform perturbations do, despite maintaining integrability, as indicated by the Poisson distribution. In stark contrast, class (iii) perturbations induce Wigner-Dyson statistics even for infinitesimal strength, signaling the onset of quantum chaos.
Eigenstate thermalization.– The emergence of chaos in long-range systems as the interaction exponent $\alpha$ is tuned from zero to small non-zero values was previously reported in Ref. [56]. Despite the Wigner-Dyson level spacing distribution, which is a hallmark of quantum chaos, that work along with Ref. [57] suggested that the eigenstates may fail to satisfy ETH. More recently, this violation was reinforced from an open quantum system perspective in [62], and efforts to restore thermalization for long-range systems have also been proposed [63]. In this work, we revisit the claim of ETH breakdown in quantum systems with strong long-range interactions and, in contrast, provide evidence that ETH does hold, although in a more subtle, sector-dependent form.
Figure 2: Verification of the eigenstate thermalization hypothesis (ETH) for Hamiltonian (1) with $\alpha=10^{-4}$, $L=14$. The top panels show results for the full energy spectrum, with the vertical red line marking the most populated band analyzed in the bottom panels. (a) and (d): Density of states, (b) and (e): Entanglement entropy, (c) and (f): Eigenstate expectation values of $\hat{S}_{z}$, and (g): Off-diagonal elements of $\hat{S}_{z}$ for 200 eigenstates in the middle of the selected energy band. (f) and (g): Diagonal and off-diagonal ETH, respectively, are satisfied. Parity and inversion symmetries are taken into account.
For an infinitesimally small $\alpha$, the band structure of the DOS remains intact [Fig. 2(a)], although the degeneracies within each band are lifted. In the weak perturbation regime, the bands can still be approximately characterized by the total spin quantum number, which effectively acts as a quasi-conserved quantity. As a result, a proper analysis of ETH must be performed within individual quasi-sectors defined by this emergent symmetry. Constructing energy shells that mix states across different quasi-sectors can obscure the existence of chaotic states and misleadingly indicate a breakdown of ETH. This is indeed suggested by the fragmented results for the scaled half-chain entanglement entropy,
$\tilde{S}^{\mathrm{ent}}_{L/2}=-\frac{2}{L\ln 2}\text{Tr}_{L/2}\left[\hat{\rho}_{L/2}\ln{(\hat{\rho}_{L/2})}\right]$ (3)
in Fig. 2(b), and the eigenstate expectation values, $\langle n|\hat{S}_{z}|n\rangle$, of the total magnetization in the $z$-direction in Fig. 2(c), both shown as a function of the energy levels of the entire spectrum.
This seemingly violation of ETH disappears when the analysis is restricted to individual energy bands. As shown in Fig. 2(d), the DOS for the most populated band closely resembles a Gaussian distribution, as typical of many-body quantum systems [64]. Consistent with the presence of chaotic eigenstates [65, 49], which underlie the validity of ETH [66], the entanglement entropy of the eigenstates within the band, and away from its edges, exhibits a smooth dependence on energy in Fig. 2(e). This behavior is mirrored by the smooth variation of the eigenstate expectation values of $\hat{S}_{z}$, as shown in Fig. 2(f). Additionally, the distribution of the off-diagonal matrix elements of $\hat{S}_{z}$ for the eigenstates in the middle of the selected energy band, displayed in Fig. 2(g), is well-approximated by a Gaussian, in agreement with the predictions of off-diagonal ETH [67, 68].
These observations provide strong evidence that ETH is satisfied within individual energy bands. Furthermore, we have confirmed that this conclusion holds robustly for other perturbations within class (iii).
The onset of chaos for arbitrarily small perturbations was recently reported also in [69] in similar scenarios of highly degenerate energy levels. However, contrary to our analysis and that in Ref. [56], the unperturbed Hamiltonians in [69] are non-interacting and the perturbations involve nearest-neighbor interactions.
Infinitesimal perturbation.– To understand why certain perturbations lead to chaos while others preserve integrability in the long-range model, we employ degenerate perturbation theory around the analytically tractable limit $\alpha=0$. Introducing a perturbation $\hat{V}$ from one of the three aforementioned classes, the perturbed Hamiltonian takes the form
$\hat{H}_{\mathrm{perturbed}}=\hat{H}^{\alpha=0}+\hat{V}.$ (4)
We focus on the most populated degenerate energy band of the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$, with eigenstates $|n_{0}\rangle$, and complement our analysis of the eigenvalues of the perturbed Hamiltonian in Eq. (4) with that of the first-order energy corrections. To this end, we construct the matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ within the selected band and diagonalize it to obtain the energy corrections $\lambda_{n_{0}}$. If the eigenvalues $\lambda_{n_{0}}$ are all zero or show internal degeneracies, the original degeneracy of the band is preserved or partially lifted, suggesting that integrability is maintained. In contrast, a fully non-degenerate $\lambda_{n_{0}}$ accompanied by level repulsion signals the breakdown of integrability and the onset of quantum chaos. Degenerate perturbation theory thus provides a diagnostic of the interplay between symmetry, degeneracy lifting, and the onset of chaos, as shown next.
Class (i): A longitudinal one-body impurity, such as $\hat{\sigma}^{x}_{L/2}$, leaves $\lambda_{n_{0}}=0$, implying degeneracy and persistent integrability. A transverse impurity $\hat{\sigma}^{z}_{L/2}$ lifts the degeneracy but induces a structured band splitting with residual degeneracies, again consistent with integrability. Similar features arise for two-body local perturbations, such as $\hat{\sigma}^{x}_{L/2-1}\hat{\sigma}^{x}_{L/2}$, where degeneracies are lifted in a structured manner, resulting in additional subbands.
Class (ii): Homogeneous fields, such as $\sum_{i}\hat{\sigma}^{x}_{i}$, preserve the global $SU(2)$ symmetry and hence integrability. The situation becomes more intricate with inhomogeneous fields. Transverse inhomogeneous fields, for instance $\sum_{i}h_{i}\hat{\sigma}^{z}_{i}$, break symmetry and lift degeneracies, yet for either weak ($\delta\ll J$) or strong ($\delta\sim{\cal O}(J)$) disorder, we observe Poissonian level statistics for $\lambda_{n_{0}}$, indicating preserved integrability. However, longitudinal inhomogeneous fields, $\sum_{i}h_{i}\hat{\sigma}^{x}_{i}$, lead to level statistics for $\lambda_{n_{0}}$ intermediate regime between Poisson and Wigner-Dyson distributions, suggesting level repulsion in the presence of possible (quasi-)symmetries.
To further assess level repulsion and ETH in this setting, we return to the exact eigenvalues of the Hamiltonian in Eq. (4). We consider very weak random perturbations ($\delta=10^{-4}$) and analyze in Fig. 3 different indicators of quantum chaos for the most populated energy band. The mean ratio of consecutive level spacings, $\bar{r}$, exhibits strong fluctuations across individual disorder realizations, as indicated by the relatively broad distributions in Fig. 3(a). However, the ensemble average, $\langle\bar{r}\rangle$, increases with system size $L$, as evidenced in Fig. 3(a) by a systematic shift toward larger values and narrowing of the distributions as $L$ increases. The inset confirms this trend, showing that $\langle\bar{r}\rangle$ grows with $L$.
In Fig. 3(b) we examine the spectral form factor, defined as [59]
$K(t)=\left|\frac{1}{\mathcal{D}}\sum_{n=1}^{\mathcal{D}}e^{iE_{n}t}\right|^{2},$ (5)
where $\mathcal{D}$ is the Hilbert space dimension of the most populated energy band. This quantity sensitively captures spectral correlations by developing a dip–ramp–plateau structure (correlation hole) even in the presence of symmetries [54, 70]. The correlation hole, corresponding to the time interval with values of $K(t)$ below its saturation (dashed) line, is seen in Fig. 3(b) for all three system sizes considered, with no sign of reduction of the relative depth with $L$.
Figures 3(a)-(b) suggest that integrability is broken for arbitrarily weak random longitudinal fields, yet quantum chaos is not fully developed. This is further supported by Fig. 3(c), where off-diagonal ETH is probed with the distribution of the matrix elements of $\hat{S}_{z}$, as in Fig. 2(g). Instead of the Gaussian form expected in many-body quantum chaos, we observe a log-normal distribution, indicative of structured eigenstates potentially shaped by residual (quasi-)symmetries. These results point to the random longitudinal field perturbation as a subtle and nontrivial case deserving further theoretical and numerical exploration.
Figure 3: Analysis of (a)-(b) spectral correlations and (c) off-diagonal ETH for the most populated energy band of the perturbed Hamiltonian in Eq. (4) with $\hat{V}=h_{i}\sigma_{i}^{x}$ , random numbers $h_{i}\in[-\delta,\delta]$ and $\delta=10^{-4}$. (a) Distribution of the mean value of the ratio of consecutive level spacings, $\bar{r}$, for various disorder realizations and their average vs system size $L$ in the inset. (b) Spectral form factor for different system sizes; dashed line indicates the saturation value. (c) Distribution of the off-diagonal elements of $\hat{S}_{z}$ for 200 levels in the middle of the energy band, $L=13$. The solid line represents a log-normal distribution.
Class (iii): These perturbations break the $SU(2)$ symmetry and immediately lift all degeneracies within the originally degenerate energy band. The resulting first-order energy corrections exhibit Wigner-Dyson statistics, even for tiny perturbation strengths, unambiguously signaling the onset of many-body quantum chaos. The emergence of level repulsion already at the level of first-order perturbation theory reveals the microscopic mechanism by which class (iii) perturbations induce chaos. These observations are consistent across various symmetry-resolved sectors and are robust to changes in the specific form of the extensive interaction.
Conclusion.– We have shown that a fully connected many-body quantum system can display striking resilience to certain classes of perturbations, while others trigger quantum chaos even at infinitesimally small strengths. Drawing a heuristic analogy with classical chaos, where localized chaotic regions in phase space emerge and progressively expand until they merge together, quantum chaos in fully connected systems unfolds in a similarly fragmented fashion [71]. It first emerges within individual energy bands associated with quasi-conserved quantities, and as the strength of the integrability-breaking perturbation increases, these chaotic regions broaden and eventually coalesce, signaling a global breakdown of integrability.
A natural theoretical and experimental extension of this work is to investigate the nonequilibrium dynamics and thermalization timescales near $\alpha=0$, comparing the behavior across different perturbation classes. In particular, it is important to contrast the evolution of initial states confined to a single energy band with that of states that span multiple bands.
Acknowledgements.– The authors thank Shamik Gupta for inspiring discussions on long-range systems. S. K. P. acknowledges helpful discussions with Marcello Dalmonte, Tobias Schatez, Shreya Vardhan, and Vyshakh B. R.  L. F. S. thanks start-up funding from the University of Connecticut. S. K .P. is supported at the Tata Institute of Fundamental Research (TIFR) through a graduate fellowship from the Department of Atomic Energy (DAE), India. The authors gratefully acknowledge the “School on Classical and Quantum Long-range Interacting Systems 2024” held at TIFR Mumbai, where this project was initiated.
Supplemental Material:
Fragmented eigenstate thermalization versus robust integrability in long-range models
Soumya Kanti Pal1, Lea F. Santos2
1Department of Theoretical Physics, Tata Institute of Fundamental Research, Homi Bhabha Road, Mumbai 400005, India
2Department of Physics, University of Connecticut, Storrs, Connecticut 06269, USA
This supplemental material provides additional figures and discussions that support the findings of the main text. It is organized into three parts: (a) we first explain how introducing a small $\alpha$ can be interpreted as a perturbation to the Hamiltonian at $\alpha=0$; (b) we then describe the band structure of the $\alpha=0$ Hamiltonian [Eq. (2) in the main text] and its characterization; and (c) finally, we present the degenerate perturbation scheme within the most populated band at $\alpha=0$, detailing the first-order corrections for various perturbation classes and how they capture the emergence of integrability or the onset of chaos.
I Justification on why tuning $\alpha$ is a non-local perturbation to $\hat{H}^{\alpha=0}$
s Starting from the Hamiltonian in Eq. (1) of the main text, we consider a small deviation $\alpha\ll 1$ from $\alpha=0$. The Hamiltonian can then be expressed as
$\hat{H}^{\alpha\ll 1}=\hat{H}^{\alpha=0}+\underbrace{(\hat{H}^{\alpha\ll 1}-\hat{H}^{\alpha=0})}_{\equiv\,\hat{V}_{\alpha}},$ (S6)
where the second term, $\hat{V}_{\alpha}$, acts as a non-local perturbation. To quantify the perturbative nature of $\hat{V}_{\alpha}$ for $\alpha<1$, we define
$\epsilon(\alpha)=\frac{||\hat{V}_{\alpha}||}{||\hat{H}^{\alpha=0}||},$ (S7)
where the Hilbert-Schmidt norm is given by $||A||=\sqrt{\sum_{i}\lambda_{i}^{2}}$, with $\lambda_{i}$ denoting the eigenvalues of the Hermitian matrix $A$. As shown in Fig. S4, we find $\epsilon(\alpha)<1$ for $\alpha<1$, validating the interpretation of $\hat{V}_{\alpha}$ as a perturbation to $\hat{H}^{\alpha=0}$.
Figure S4: For the Hamiltonian given in Eq. (1) of the main text with $\alpha=0$, we demonstrate that introducing a small but finite $\alpha$ acts as a perturbation, as the ratio of the norms can be fitted to the form $\epsilon=0.49\,\alpha^{1.0}$.
II Discussion on the total spin conservation at $\alpha=0$: Band structures in DOS
We begin by recalling Eq. (2) of the main text, which corresponds to the $\alpha=0$ limit of the Hamiltonian in Eq. (1):
$\hat{H}^{\alpha=0}=\frac{2J}{N}\hat{S}_{x}^{2}+2h\hat{S}_{z}-2J,$ (S8)
[…]
$\sum_{s=0}^{L/2}(2s+1)=\left(\frac{L}{2}+1\right)^{2}.$ (S9)
[…]
$n(s,L)=\frac{2s+1}{L+1}\binom{L+1}{\frac{L}{2}-s},$ (S10)
[…]
Figure S5: For the fully connected limit of the LRTFIM with the Hamiltonian given in Eq. (S8) for $L=8$ spins, panel (a) shows the density of states (DOS), revealing 25 distinct energy bands. Panel (b) confirms that $\hat{S}^{2}$ is diagonal in the computational eigenbasis of Eq. (S8), and panel (c) illustrates how each band can be uniquely identified by its corresponding total spin quantum number $s$.
Another important aspect is to determine for a given $L$, if one can find the maximum $s$-sector. This boils down to the task of maximizing the function $n(s,L)$ with respect to $s$ for a given $L$. This can be easily numerically settled by just observing that the maximal sector (the sector that has most number of eigenstates in each band) $s_{\mathrm{max}}\equiv\max_{s}{n(s,L)}$ varies as a function of $L$, as demonstrated in Fig. S6. Identifying the $s_{\mathrm{max}}$ is particularly relevant as one is usually interested in understanding ETH or thermalization dynamics in the most populated symmetry sector.
Figure S6: For the Hamiltonian at $\alpha=0$ in Eq. (S8), we present the characterization of maximally populated $s_{\mathrm{max}}$-sector for increasing $L$. Even for numerically accessible system sizes, we find that the variation of $s_{\mathrm{max}}$ with $L$ is quite relevant.
III Degenerate perturbation theory: Why a certain class gives chaos while others retain integrability?
In this section, we provide a first-principles explanation for the emergence of either integrability or chaos in the cases of class (i), class (ii), and class (iii) perturbations. To this end, we employ degenerate perturbation theory with the setup
$\hat{H}=\hat{H}^{\alpha=0}+\hat{V},$ (S11)
where $\hat{H}^{\alpha=0}$ is the unperturbed Hamiltonian as given in Eq. (2) of the main text, and $\hat{V}$ denotes the perturbation belonging to one of the classes discussed previously. As outlined earlier, the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$, owing to its $SU(2)$ symmetry, exhibits a highly degenerate density of states (DOS), organized into bands labeled by the total spin quantum number $s$. The method for analyzing the effect of perturbations is as follows:
  • First, we pin down the symmetries of the Hamiltonian $\hat{H}$ (perturbation included) and we numerically obtain the spectrum of the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$ in a symmetry-resolved manner.
  • Next, we numerically identify the most populated energy band in the DOS of $\hat{H}^{\alpha=0}$, with energy $E_{0}$, and restrict our analysis to the corresponding eigenstates $\{|m_{0}\rangle\}$ within this band.
  • For each pair of states $|m_{0}\rangle$, $|n_{0}\rangle$ in this band, we compute the matrix elements of the perturbation: $\langle m_{0}|\hat{V}|n_{0}\rangle$.
  • We diagonalize this perturbation matrix. The resulting eigenvalues $\{\lambda_{n_{0}}^{i}\}$ yield the first-order energy corrections as $E_{n}^{i}=E_{n_{0}}+\lambda_{n_{0}}^{i}$.
If most of the eigenvalues $\lambda_{i}$ vanish, the degeneracy is largely preserved, indicating integrability. If the eigenvalues are non-zero but exhibit degeneracies or clustering, the original band splits into multiple subbands, and the system retains integrable features. On the other hand, if the degeneracy is completely lifted, one must analyze the statistics of the corrected energy levels: the presence of level repulsion indicates the emergence of chaos, while level attraction or clustering suggests persistent integrability.
To proceed, we consider the three classes separately.
III.A Class (i): Non-extensive
For this class, we consider perturbations—either one-body or two-body—whose spatial support remains fixed as the system size increases.
One-body perturbations. We introduce a single-site impurity at a bulk site, either in the longitudinal or transverse direction. For a longitudinal impurity, the perturbation takes the form $\delta\hat{\sigma}^{x}_{L/2}$. In this case, we find that all first-order corrections vanish, i.e., $\lambda_{n_{0}}^{i}=0$ for all $i$, indicating that the degeneracy of the unperturbed band remains completely intact. In contrast, a transverse impurity, given by $\delta\hat{\sigma}^{z}_{L/2}$, lifts the degeneracy: all eigenvalues $\lambda_{i}$ become nonzero. However, the spectrum of the perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ exhibits internal degeneracies, leading to a splitting of the original degenerate band into several subbands, each with its own residual degeneracy.
Although perturbation theory is strictly valid only in the $\delta\ll 1$ regime, we observe that this qualitative structure persists even for larger values of $\delta$. These results explain the persistence of integrability in the presence of such local perturbations. We also note that in the other integrable limit $\alpha\to\infty$, a longitudinal impurity of the same kind would instead induce chaos, in stark contrast to the behavior observed here.
Two-body perturbations. In the bulk, we consider a single two-body interaction term, either in the longitudinal or transverse direction, given respectively by $\hat{\sigma}^{x}_{L/2-1}\hat{\sigma}^{x}_{L/2}$ or $\hat{\sigma}^{z}_{L/2-1}\hat{\sigma}^{z}_{L/2}$. Similar to the one-body case, the eigenvalues $\{\lambda_{n_{0}}^{i}\}$ of the perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ exhibit degeneracies, indicating that a single band splits into multiple subbands, each retaining some degeneracy. This spectral structure reflects the persistence of integrability despite the perturbation. We also note that the choice of the specific pair $(L/2-1,L/2)$ is not essential; any arbitrary pair in the bulk produces qualitatively similar results.
III.B Class (ii): Extensive one-body
This class of perturbations includes both homogeneous and inhomogeneous field terms in both directions, applied to all sites.
For homogeneous perturbations as in the form of $\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}$ or $\sum_{i=1}^{L}\hat{\sigma}_{i}^{z}$, the perturbed Hamiltonian still retains the original $SU(2)$ symmetry by conserving the total spin $\hat{S}^{2}$. This leads to the organization of the spectrum into more degenerate bands, thereby trivially extending the integrability of the unperturbed system.
For inhomogeneous perturbations in the transverse direction, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{z}$, with $h_{i}\in[-\delta,\delta]$ being random, we find that the degeneracies are getting lifted inside a band. Also, the level-ratio test reveals a clear Poissonian signature, implying integrability. Although the original $SU(2)$ conservation is lost, the first-order corrections reveal zero correlation, which we understand as a hint towards integrability for exact energy statistics. For random field perturbations in the longitudinal direction $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}$, the situation is trickier, and a discussion on its exact eigenvalue statistics, as well as the behaviour of the spectral form factor, and matrix elements inside a band for infinitesimal perturbation is presented in the main text.
III.C Class (iii): Extensive two-body
In this class of perturbations, we have nearest-neighbour (nn) and interactions beyond nn, evidently incorporating the perturbation in $\alpha$ by turning on infinitesimal $\alpha$ from 0.
Let us first consider the case when the perturbation is nearest-neighbor and given by
$\hat{V}^{\text{nn}}=\delta\sum_{k=1}^{L-1}\hat{\sigma}_{k}^{x}\hat{\sigma}_{k+1}^{x}.$
We begin by examining the level statistics in the central one-third of the spectrum, resolved by both parity and inversion symmetries. Even for very small values of $\delta$, we observe Wigner–Dyson statistics, indicative of quantum chaos.
To further investigate the mechanism underlying this chaotic behavior, we construct the matrix $\langle m^{0}|\hat{V}^{\text{nn}}|n^{0}\rangle$ within the most populated degenerate band of $\hat{H}^{\alpha=0}$, restricted to the relevant symmetry sector. Diagonalizing this matrix, we obtain the first-order energy corrections $\{\lambda_{n_{0}}^{i}\}$. All the eigenvalues are found to be distinct, indicating that the degeneracy is completely lifted at first order in perturbation theory.
Remarkably, even the set of first-order corrections $\{\lambda_{n_{0}}^{i}\}$ exhibits level repulsion: the distribution of the unfolded consecutive level spacings $\delta\lambda_{n}^{i}=\lambda_{{n+1}_{0}}^{i}-\lambda_{n_{0}}^{i}$, after sorting the eigenvalues energetically, follows Wigner’s surmise. The emergence of level repulsion (see Fig. S7) already at the level of the first-order corrections provides a clear mechanism for the onset of chaos in the perturbed system.
Figure S7: For the perturbation $\delta\sum_{k=1}^{L-1}\hat{\sigma}_{k}^{x}\hat{\sigma}_{k+1}^{x}$ with $\delta=10^{-4}$, in panel (a) we show the density of states with red-dashed line indicating one of the most populated bands for size $L=14$. In panel (b), the distribution of consecutive spacings amongst the first-order corrections in the band is shown to follow Wigner-Dyson statistics, underpinning the mechanism behind the onset of chaos at infinitesimal perturbation strength.
We note here that the same is true for all perturbations in this class.